5. The US DP growth rate for 2007-2017 is reported on the following chart. Determine the consecutive three year period between 2007 and 2017 with the largest growth rate Develop a model to predict the US DP growth rate for December 31, 2020 Value Date Dec 31, 2017 4.49% Dec 31, 2016 3.40% Dec 31 , 2015 2.89% Dec 31, 2014 4.42% Dec 31, 2013 4.43% Dec 31, 2012 3.56% Dec 31, 2011 3.65% Dec 31, 2010 4.19% Dec 31, 2009 0.47% Dec 31, 2008-0.83%
Answer #1

As shown in the table given in the question, the consecutive three years between 2007 and 2017 with the largest growth rate is:

Year Growth rate
31-Dec-12 3.56%
31-Dec-13 4.43%
31-Dec-14 4.42%

Now, develop a regression model in which the growth rate is function of time.

Assume t=0 for Period 31-Dec-2007

Thus, the required table is:

Year t Growth rate
31-Dec-07 0 4.59%
31-Dec-08 1 -0.83%
31-Dec-09 2 0.47%
31-Dec-10 3 4.19%
31-Dec-11 4 3.65%
31-Dec-12 5 3.56%
31-Dec-13 6 4.43%
31-Dec-14 7 4.42%
31-Dec-15 8 2.89%
31-Dec-16 9 3.40%
31-Dec-17 10 4.49%
31-Dec-18 11
31-Dec-19 12
31-Dec-20 13

The regression model is:

SUMMARY OUTPUT Regression Statistics MultipleR R Square Adjusted R Square Standard Error Observations 0.421748173 0.177871521 0.086523913 0.017027351 ANOVA MS Significance F Regression Residual Total 10.000564551 0.0005645511.947194063 0.196359964 9 0.002609376 0.000289931 10 0.003173927 Coefficients Standard Errort 0.0207272730.0096047212.158029639 0.059257222 -0.001000116 0.042454661-0.0010001160.042454661 0.002265455 0.001623494 1.3954189560.196359964 -0.001407144 0.005938054-0.001407144 0.005938054 P-value Lower 95% Up er95% Lower 95.0% Upper 95.0% InterceptThus the model is:

Growth_rate = 0.020727273 + 0.002265455 * t

For Dec 31, 2020, the value of t=13

Hence, the predicted growth rate in Dec 31, 2020 is:

= 0.020727273 + 0.002265455 * t

=0.020727273 + 0.002265455 * 13

=0.050178182

=5.0178%